a patient in a hospital is often administered an intravenous (IV) drip contains an aqueous solution, this aqueous solution contains 0.85% (mass by volume) of sodium chloride or 5% (mass by volume) of glucose. calculate the molarity of both these solution.


1
Expert's answer
2021-06-23T05:16:33-0400

This question is answered by using the simple concept of calculation of molarity of the solution using the simple formula and the mass/volume concentration of the solution.

Mass of NaCl =0.85g

MM=58.5g/mol

Volume =100mL

So, NaCl="\frac{Mass×1000}{MM×Volume(mL)}"


"=\frac{0.85×1000}{58.5×100}"


"NaCl=0.1453M"

5%(mass by volume) glucose

Mass=5g

Volume=100mL

MM=180g/mol

"Glucose =\frac{Mass×1000}{MM×Volume(mL)}"


"=\frac{5×1000}{180×100}"


"Glucose =0.2778M"